# The Meta-Representation Hypothesis

### v3

@@ 4 Central Hypothesis
The figure illustrates two randomly initialized policies independently trained using reinforcement learning algorithms. In this case, since the training samples only include a portion of the MDPs, the policies are likely to overfit to irrelevant features and fail to converge to a robust hypothesis space.
Figure 3: Generalization performance from 500 levels in Procgen benchmark with different methods. The mean and standard deviation are shown across 3 seeds. Our proposed DML-based method gains significant improvement compared with the baseline algorithm.
@@
Introducing the DML loss into the training process of two policies (denoted as policy A and policy B) facilitates mutual learning, which can mitigate overfitting to irrelevant features. Due to the random initialization of policies A and B, they generate different training samples. The DML loss encourages both policies to make consistent decisions on the same observations. As a result, any irrelevant features learned by policy A are likely to degrade the performance of policy B (see Appendix B for further explanation), and vice versa. As training progresses, DML will drive both policies to learn more meaningful and useful representations, gradually reducing the divergence between them (right of the Figure 2). Ideally, we hypothesize that both policies will converge to meta-representations that capture the essential aspects of high-dimensional observations as time grows.
An intriguing analogy for our hypothesis is the process of truth emergence. Typically, each scholar offers their unique perspective, but for it to be widely accepted, it must garner consensus from peers within the field, or even from the broader academic community. We can draw a parallel between DML and the peer review process: when a particular viewpoint is accepted by the majority, it is more likely to reflect an objective truth—though, of course, this does not preclude the possibility that everyone could be mistaken, as seen during the era when geocentrism was widely endorsed. On a deeper level, our hypothesis aligns with the philosophical concept of convergent realism (Laudan, 1981; Kelly & Glymour, 1989; Huh et al., 2024), which posits that science progresses towards an objective truth.
@@ 5.1 Implementation Details
We use Procgen (Cobbe et al., 2019, 2020) as the experimental benchmark for testing generalization performance. Procgen is a suite of 16 procedurally generated game-like environments designed to benchmark both sample efficiency and generalization in reinforcement learning, and it has been widely used to test the generalization performance of various reinforcement learning algorithms (Wang et al., 2020; Raileanu & Fergus, 2021; Raileanu et al., 2021; Lyle et al., 2022; Rahman & Xue, 2023; Jesson & Jiang, 2024).
We employ the Proximal Policy Optimization (PPO) (Schulman et al., 2017; Cobbe et al., 2020) algorithm as our baseline, as PPO is one of the most widely used model-free reinforcement learning algorithms. Specifically, given a parameterized policy $\pi_{\theta}$ ($\theta$ represents the parameters), the objective of $\pi_{\theta}$ is to maximize
$\mathop{\mathbb{E}}_{(o_{t},a_{t})\sim\pi_{\theta_{\mathrm{old}}}}\left\{\min\left[r_{t}(\theta)\cdot\hat{A}(o_{t},a_{t}),\tilde{r}_{t}(\theta)\cdot\hat{A}(o_{t},a_{t})\right]\right\},$ (19)
where $\hat{A}$ is the advantage estimate, and
$r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}|o_{t})}{\pi_{\theta_{\mathrm{old}}}(a_{t}|o_{t})},\enspace\tilde{r}_{t}(\theta)=\mathrm{clip}\left(r_{t}(\theta),1-\epsilon,1+\epsilon\right),$ (20)
with $\pi_{\theta_{\mathrm{old}}}$ and $\pi_{\theta}$ being the old policy and the current policy.
We randomly initialize two agents to interact with the environment and collect data separately. Similar to the DML loss (5) used in supervised learning, we also introduce an additional KL divergence loss term, which is
@@
where $\mathcal{L}_{\mathrm{RL}}$ is the reinforcement learning loss and $\mathcal{L}_{\mathrm{KL}}$ is the KL divergence loss, $\alpha$ is the weight. And then we optimize the total loss of both agents, which is the average of their DML losses, as shown in Algorithm 2.
Finally, we do not claim to achieve state-of-the-art (SOTA) performance, but rather to verify that the DML technique indeed helps agents learn more robust representations from high-dimensional observations and leads to consistent improvements in generalization performance, providing empirical support for our central theory and hypothesis.
Algorithm 2 PPO with DML
@@
Figure 4: To test the robustness of the trained policy, we obfuscate the agent’s observations using convolutional layers randomly initialized with a standard Gaussian distribution. If the agent has indeed learned to ignore irrelevant features from noisy observations, it should exhibit better robustness to such obfuscations. Notably, the feature extraction of the PPO encoder enhanced by DML is highly stable and focused (red points), whereas the features extracted by the original PPO encoder are significantly dispersed (blue points).
@@
PPO $6.15\pm 1.58$ $8.19\pm 0.96$ $8.60\pm 0.73$ $14.40\pm 1.19$ $0.62\pm 0.48$ $1.70\pm 0.35$ $0.09\pm 0.05$ $6.41\pm 1.25$
PPO with DML $3.91\pm 0.58$ $0.32\pm 0.20$ $1.38\pm 0.35$ $4.66\pm 0.70$ $0.09\pm 0.06$ $1.57\pm 0.41$ $1.29\pm 0.16$ $1.22\pm 0.63$
@@
PPO $1.38\pm 0.19$ $10.92\pm 1.80$ $4.94\pm 1.57$ $5.79\pm 1.10$ $12.44\pm 3.28$ $5.91\pm 1.30$ $4.61\pm 1.00$ $3.72\pm 0.93$
PPO with DML $0.05\pm 0.03$ $0.75\pm 0.29$ $2.33\pm 1.07$ $1.44\pm 0.33$ $2.08\pm 0.86$ $2.30\pm 1.07$ $2.44\pm 0.36$ $2.88\pm 0.75$
Figure 5: Generalization performance of retraining policies using the frozen encoders obtained from the PPO baseline and our method.
5.2 Empirical Results
We compare the generalization performance of our approach against the PPO baseline on the Procgen benchmark, under the hard-level settings (Cobbe et al., 2020). The results are illustrated in Figure 3. It can be observed that DML technique indeed leads to consistent improvements in generalization performance across all environments. Notably, for the bigfish, dodgeball, and fruitbot environments, we have observed significant improvements. Moreover, the experimental results for all 16 environments in Procgen benchmark, including training performance and generalization performance, can be found in Appendix C.
A natural concern arises: how can we determine whether DML improves generalization performance by enhancing the policy robustness against irrelevant features, or simply due to the additional information sharing between these two agents during training (each agent receives additional information than it would from training alone)? To answer this question, we conducted robustness testing on the trained policies in Section 5.3 and added an ablation study in Section 5.4 to verify our theory and hypothesis.
5.3 Robustness Testing
To further verify that our method has indeed learned more robust policies, we design a novel approach to test policy robustness against irrelevant features, as shown in Figure 4. For each current frame, we input it into multiple convolutional layers randomly initialized with a standard Gaussian distribution, and then compute the average KL divergence of the policy before and after the perturbation by these random convolutional layers. This allows us to effectively test the robustness of the trained policies without changing the underlying semantics. The results can be seen from Table 1. We can see that the average changes in KL divergence of our method is lower than the PPO baseline across almost all environments, with a smaller standard deviation, providing strong empirical support for our central hypothesis.
Moreover, we employ t-SNE to visualize the agent’s encoding of high-dimensional observations in the bigfish environment, as shown in Figure 4. Each scatter point represents a low-dimensional embedding of a vector obtained by passing the current pixel input through these random convolutional layers, and then fed into the agent’s encoder. It can be observed that the scatter points of our method are more tightly clustered, indicating a more robust representation of high-dimensional noisy observations, which serves as further strong evidence for our hypothesis.
5.4 Ablation Study
To verify that the generalization performance of the agent benefits from more robust policies, we designed additional ablation experiments. Specifically, we used the frozen encoders obtained from the PPO baseline and our method to retrain the policies, the results are shown in Figure 5. Since the policy obtained from our method is more robust to irrelevant features (as demonstrated in Section 5.3), the encoder learns a better representation of the high-dimensional observations. Therefore, based on our theoretical results, retraining policies using the frozen encoders obtained from our method should have better generalization performance. We can see that the generalization performance in Figure 5 strongly supports our theoretical results.
In summary, Section 5.2 validates the effectiveness of DML technique for generalization, Section 5.3 verifies our central hypothesis, and Section 5.4 confirms our theoretical results.
6 Related Work
The generalization of deep reinforcement learning has been widely studied, and previous work has pointed out the overfitting problem in deep reinforcement learning (Rajeswaran et al., 2017; Zhang et al., 2018a; Justesen et al., 2018; Packer et al., 2018; Song et al., 2019; Cobbe et al., 2019; Grigsby & Qi, 2020; Cobbe et al., 2020; Yuan et al., 2024).
A natural approach to avoid the overfitting problem in deep reinforcement learning is to apply regularization techniques originally developed for supervised learning such as dropout (Srivastava et al., 2014; Farebrother et al., 2018; Igl et al., 2019), data augmentation (Laskin et al., 2020; Kostrikov et al., 2020; Zhang & Guo, 2021; Raileanu et al., 2021; Ma et al., 2022), domain randomization (Tobin et al., 2017; Yue et al., 2019; Slaoui et al., 2019; Mehta et al., 2020), or network randomization technique (Lee et al., 2019).
On the other hand, in order to improve sample efficiency, previous studies encouraged the policy network and value network to share parameters (Schulman et al., 2017; Huang et al., 2022). However, recent works have explored the idea of decoupling the two and proposed additional distillation strategies (Cobbe et al., 2021; Raileanu & Fergus, 2021; Moon et al., 2022). In particular, Raileanu & Fergus (2021) demonstrated that more information is needed to accurately estimate the value function, which can lead to overfitting.
7 Conclusion
In this paper, we provide a novel theoretical framework to explain the generalization problem in deep reinforcement learning. We also hypothesize that the DML technique facilitates meta-representation learning. Strong empirical results support our central theory and hypothesis, demonstrating that our approach can improve the generalization performance of RL systems by enhancing robustness against irrelevant features. Our work provides valuable insights into the development of more adaptable and robust RL systems capable of generalizing across diverse domains.
8 Discussion
A key insight from our work is the process of extracting patterns from empirical observations, a powerful abstraction ability that is central to human cognition. This raises a fundamental question: if human perception is based on electrical and chemical signals in the brain, then how can we infer the true nature of the world?
@@ 8 Discussion
Our approach offers a potential answer through the concept of cognitive alignment (Falandays & Smaldino, 2022). By encouraging agents to make consistent decisions based on the same observations, our method fosters a process akin to cognitive alignment, which has been fundamental in human societal development. For instance, in voting, the majority rule is employed because decisions supported by the majority are perceived as more reliable. Similarly, our method facilitates cognitive alignment between agents, enabling them to converge on objective truths despite noisy or irrelevant features in their observations.
Over time, the cognitive alignment between agents encourages the convergence of their individual representations toward a more accurate understanding of the environment. This process mirrors the philosophical notion presented in Plato’s Allegory of the Cave (Cohen, 2006), where individuals break free from the constraints of their limited perceptions to grasp the true nature of reality.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Proofs

### v4

@@ 4.2 Central Hypothesis
Motivated by Section 4.1, DML can be viewed as a form of implicit regularization against irrelevant features, as demonstrated in Figure 2, which illustrates two randomly initialized policies independently trained using reinforcement learning algorithms. In this case, since the training samples only include a portion of all possible MDPs, the policies are likely to overfit to irrelevant features and fail to converge to a robust hypothesis space.
@@
Applying DML to the training process of both policies facilitates mutual learning, which can mitigate overfitting to irrelevant features. Due to the randomness of parameter initialization and the interaction process, they generate different training samples, DML encourages both policies to make consistent decisions based on the same observations. As discussed in Section 4.1, any irrelevant features learned by policy A are likely to degrade the performance of policy B, and vice versa. As training progresses, DML will drive both policies to learn more meaningful and useful representations, gradually reducing the divergence between them. Ideally, we hypothesize that both policies will capture the essential aspects of high-dimensional observations as time grows.
@@ 5 Experiments
This section presents our main empirical results. Section 5.1 introduces the implementation details, Section 5.2 validates the effectiveness of DML technique for improving generalization performance, Section 5.3 verifies our central hypothesis, and Section 5.4 confirms our theoretical results.
@@ 5.1 Implementation Details
We use Procgen [4, 5] as the experimental benchmark for testing generalization performance. Procgen is a suite of 16 procedurally generated game-like environments designed to benchmark both sample efficiency and generalization in reinforcement learning, and it has been widely used to test the generalization performance of various reinforcement learning algorithms [45, 33, 34, 25, 32, 18].
We employ the Proximal Policy Optimization (PPO) [37, 5] as our baseline. Specifically, given a parameterized policy $\pi_{\theta}$ ($\theta$ represents the parameters), the objective of $\pi_{\theta}$ is to maximize
$J(\theta)=\mathop{\mathbb{E}}_{(o_{t},a_{t})\sim\pi_{\theta_{\mathrm{old}}}}\left\{\min\left[r_{t}(\theta)\cdot\hat{A}(o_{t},a_{t}),\mathrm{clip}\left(r_{t}(\theta),1-\epsilon,1+\epsilon\right)\cdot\hat{A}(o_{t},a_{t})\right]\right\},$ (17)
where $\hat{A}$ is the advantage estimate, and $r_{t}(\theta)=\pi_{\theta}(a_{t}|o_{t})/\pi_{\theta_{\mathrm{old}}}(a_{t}|o_{t})$ is the probability ratio, where $\pi_{\theta_{\mathrm{old}}}$ and $\pi_{\theta}$ denote the old and current policies, respectively.
We randomly initialize two agents to interact with the environment and collect data separately. Similar to the DML loss (5) used in supervised learning, we also introduce an additional KL divergence loss term, which leads to
@@
where $\mathcal{L}_{\mathrm{RL}}$ is the reinforcement learning loss and $\mathcal{L}_{\mathrm{KL}}$ is the KL divergence loss, $\alpha$ is the weight. And then we optimize the total loss of both agents, which is the average of their DML losses, as shown in Algorithm 1, which we name Mutual Distillation Policy Optimization (MDPO).
Algorithm 1 Mutual Distillation Policy Optimization (MDPO)
@@
Ultimately, we do not claim to achieve state-of-the-art (SOTA) performance, but rather provide empirical evidence for the non-trivial insight that DML serves as an implicit regularization against irrelevant features, leading to consistent improvements in generalization performance. We also acknowledge the methodological similarities with prior work such as Zhao and Hospedales [54]; despite that, we introduce representation convergence (Section 4.2), a novel insight with further supported by strong theoretical analysis (Section 3), constituting our additional contributions.
5.2 Empirical Results
We compare the generalization performance of our MDPO against the PPO baseline on the Procgen benchmark, under the hard-level settings [5], the results are illustrated in Figure 3. It can be observed that DML technique indeed leads to consistent improvements in generalization performance across all environments. Notably, for the bigfish, dodgeball, and fruitbot environments, we have observed significant improvements. Moreover, the full experimental results for all environments, including training and generalization performance, are provided in Appendix E.
A natural concern arises: how can we determine whether DML improves generalization performance by enhancing the policy robustness against irrelevant features, or simply due to the additional information sharing between these two agents during training (each agent receives additional information than it would from training alone)? To answer this question, we conducted robustness testing in Section 5.3 and added an ablation study in Section 5.4 to support our theory and hypothesis.
Figure 3: Generalization performance from 500 levels in Procgen benchmark with different methods. The mean and standard deviation are shown across 3 seeds. Our MDPO gains significant performance improvement compared with the baseline algorithms.
Figure 4: To test the robustness of the trained policy, we obfuscate the agent’s observations using convolutional layers randomly initialized with a standard Gaussian distribution. Notably, the feature extraction of MDPO is highly stable and focused (red points), whereas the features extracted by the original PPO encoder are significantly dispersed (blue points).
@@ 5.2 Empirical Results
PPO $8.19\pm 0.96$ $8.60\pm 0.73$ $6.41\pm 1.25$ $1.38\pm 0.19$ $10.92\pm 1.80$ $5.79\pm 1.10$ $12.44\pm 3.28$
MDPO (ours) $0.32\pm 0.20$ $1.38\pm 0.35$ $1.22\pm 0.63$ $0.05\pm 0.03$ $0.75\pm 0.29$ $1.44\pm 0.33$ $2.08\pm 0.86$
5.3 Robustness Testing
To further verify that our method has indeed learned more robust policies, we design a novel approach to test policy robustness against irrelevant features, as shown in Figure 4. For each current frame, we generate adversarial samples using multiple convolutional layers randomly initialized with a standard Gaussian distribution, and record the mean and standard deviation of the KL divergence between the policy’s output for the original input and each corresponding adversarial samples. This allows us to effectively test the robustness of the trained policies without changing the underlying semantics, the results can be seen from Table 1. We can see that the average changes in KL divergence of our MDPO policy is lower than the PPO baseline across all environments, with smaller standard deviations, providing strong empirical support for our central hypothesis.
Moreover, we employ t-SNE to visualize the agent’s encoding of high-dimensional observations in the bigfish environment, as shown in Figure 4. Each point represents a low-dimensional embedding of a particular adversarial sample encoded by the agent’s encoder. It can be observed that the scatter points of our method are more tightly clustered, indicating a more robust representation of high-dimensional noisy observations, which serves as further strong evidence for our hypothesis.
5.4 Ablation Study
We design additional ablation experiments. Specifically, we use the frozen encoders obtained from the PPO baseline and MDPO to retrain the policies, and the results are shown in Figure 5. Since the policy obtained from our method is more robust (as demonstrated in Section 5.3), the encoder learns a better representation of the high-dimensional inputs. We can see that the generalization performance in Figure 5 strongly supports our theoretical results. In addition, we double the model size, batch size, and total number of interactions for the PPO baseline, as shown in Figure 6. The baseline algorithm still fails to match the performance of MDPO, demonstrating that naively scaling up the PPO baseline does not lead to stable improvements in generalization performance.
Figure 5: Generalization performance of retraining policies using the frozen encoders obtained from the PPO baseline and our proposed MDPO.
Figure 6: Generalization performance of PPO baseline with double model size, batch size, and total number of interactions, compared to original PPO and MDPO (more results in Figure 12).
6 Conclusion
In this paper, we provide a novel theoretical framework to explain the generalization problem in deep reinforcement learning. We further hypothesize that DML, as a form of implicit regularization, effectively prevents the policy from overfitting to irrelevant features. Strong empirical results support our central theory and hypothesis, demonstrating that our approach can improve the generalization performance of reinforcement learning systems by enhancing robustness against irrelevant features. Our work provides valuable insights and elegant solutions into the development of more adaptable and robust policies capable of generalizing across diverse domains.
Appendix A Related Work
The generalization of deep reinforcement learning has been widely studied, and previous work has pointed out the overfitting problem in deep reinforcement learning [35, 51, 19, 30, 41, 4, 14, 5, 49]. A natural approach to avoid the overfitting problem in deep reinforcement learning is to apply regularization techniques originally developed for supervised learning such as dropout [42, 11, 17], data augmentation [22, 48, 52, 34, 26], domain randomization [43, 50, 40, 27], or network randomization technique [24]. On the other hand, in order to improve sample efficiency, previous studies encouraged the policy network and value network to share parameters [37, 15]. However, recent works have explored the idea of decoupling the two and proposed additional distillation strategies [6, 33, 28]. In particular, Raileanu and Fergus [33] demonstrated that more information is needed to accurately estimate the value function, which can lead to overfitting.
Knowledge distillation is a learning paradigm that aims to align the student network with the teacher network to achieve knowledge transfer. A commonly used practice is to distill the knowledge learned by a large model into a smaller model to reduce inference costs after deployment [47]. On the other hand, distillation technique can also be used to distill a model with privileged information into a model with access to only partial information to improve its generalization ability. However, research has shown that knowledge distillation can also be applied to multiple student networks during training to encourage them to learn from each other, called deep mutual learning (DML) [53]. Building upon this observation, Zhao and Hospedales [54] further demonstrate that DML can improve the generalization performance of reinforcement learning agents, yet no in-depth analysis of why this happens. In addition, recent studies suggest that aligning the student networks at the output layer may be suboptimal, and recommend alignment at the logits layer instead [8, 44].
Appendix B Limitations
Since the DML loss requires explicitly calculating the KL divergence between the two policies, it is currently only suitable for policy classes that explicitly model the action distribution.
Appendix C Hyperparameters
Table 2 shows the detailed hyperparameter settings in our code, with the main hyperparameters consistent with the hard-level settings in Cobbe et al. [5], except that we train for 50M steps instead of 200M. We train the policy on the initial 500 levels and then test its generalization performance across the full distribution of levels. In addition, all experiments are conducted on a single NVIDIA A100 GPU with 80GB of memory.
Table 2: Detailed hyperparameters in Procgen.
@@ Appendix C Hyperparameters
Number of workers 64 64
Horizon 256 256
Learning rate 0.0005 0.0005
Learning rate decay No No
Optimizer Adam Adam
Total interaction steps 50M 50M
Update epochs 3 3
Mini-batches 8 8
Batch size 16384 16384
Mini-batch size 2048 2048
Discount factor $\gamma$ 0.999 0.999
GAE parameter $\lambda$ 0.95 0.95
Value loss coefficient $c_{1}$ 0.5 0.5
Entropy loss coefficient $c_{2}$ 0.01 0.01
Clipping parameter $\epsilon$ 0.2 0.2
KL divergence weight $\alpha$ 1.0
Appendix D Philosophical Insight
An intriguing analogy for our hypothesis is the process of truth emergence. Typically, each scholar offers their unique perspective, but for it to be widely accepted, it must garner consensus from peers within the field, or even from the broader academic community. We can draw a parallel between DML and the peer review process: when a particular viewpoint is accepted by the majority, it is more likely to reflect an objective truth. Going deeper, our hypothesis aligns with the philosophical concept of convergent realism [23, 21, 16], which posits that science progresses towards an objective truth.
In the Allegory of the Cave [7], Plato describes a group of people who have been chained in a cave from birth, facing the wall. They can only see the shadows cast on the wall by objects behind them, illuminated by a fire. These shadows are all they know and perceive as reality. One day, a prisoner is freed and exits the cave. Upon seeing the world outside, he discovers that the true reality is entirely different from the shadows, and he realizes that what they thought was “real” was only an illusion.
The Allegory of the Cave explores the concept of different levels of knowledge. From a metaphysical standpoint, Plato is distinguishing between the “appearance” and the “reality”. The shadows in the cave represent the sensory world, which is just an illusion, while the outside world symbolizes the realm of true reality, which can only be apprehended through rational thought and philosophical inquiry. Plato argues that what we perceive with our senses is not the true essence of reality, what he called the ideal reality, but rather a mere shadow of the higher, eternal truths.
However, even though the freed prisoner has seen the world outside the cave, how can he be certain that what he now perceives is not a shadow of something even more fundamental? This is exactly skepticism, the philosophical stance that questions the possibility of certain knowledge. It suggests that every layer of perceived reality might itself be an illusion, prompting thinkers from Descartes onward to doubt not only the senses but even the existence of the external world, demanding ever more rigorous foundations for truth. To counter skepticism, some philosophers have found it necessary to expand the definition of “reality” and propose the virtual realism [3].
Figure 7: Picasso’s The Bull [38]. By focusing on and exaggerating specific details, rather than trying to capture every detail realistically, artists can convey the core meaning or essence of the subject, more powerfully.
@@ Appendix D Philosophical Insight
This paper offers another powerful perspective against skepticism: if we regard the underlying state $s_{t}$ as the essence of things, and the rendering function $f$ as the projection (possibly a composition of multiple layers of projections), then our perception of the world becomes an observation $o_{t}=f(s_{t})$. By encouraging agents to make consistent decisions based on the same observations, our method fosters a process akin to cognitive alignment [10], which has been fundamental in human societal development. For instance, in voting, the majority rule is employed because decisions supported by the majority are perceived as more reliable. Similarly, our method facilitates cognitive alignment between agents, enabling them to converge on objective truths despite noisy or irrelevant features in their observations. Over time, the cognitive alignment between agents encourages the convergence of their individual representations toward a more accurate understanding of the environment.
It is worth noting that metaphysical ideas can also be observed in certain works of art, such as Picasso’s The Bull (Figure 7). Most people think that art is all about seeing more detail, but it is really about seeing less. Seeing basic patterns amongst the “noise”; seeing basic forms amongst the complex; seeing the few important details which convey the majority of meaning.
Appendix E More Empirical Results
E.1 Full Results
E.2 Other Baseline
We select Simple Policy Optimization (SPO) [46] as another baseline and compare the training and generalization performance of SPO with DML-based SPO, the results are shown in Figure 10 and 11.
Figure 10: Generalization performance of SPO with DML-based SPO. The mean and standard deviation are shown across 3 seeds.
Figure 11: Training performance of SPO with DML-based SPO. The mean and standard deviation are shown across 3 seeds.
We can see that DML can also significantly improve the generalization performance of SPO, which fully demonstrates that DML-based policy optimization is a general learning framework for enhancing the generalization ability of diverse reinforcement learning algorithms.
E.3 More Ablation Results
Here, we additionally present the training curves from the Ablation Study (Section 5.4), as shown in Figure 12.
Figure 12: Training performance of PPO baseline with double model size, batch size, and total number of interactions, compared to original PPO and MDPO.
Interestingly, although the scaled-up PPO nearly matches MDPO in training performance during the final stage of training in the fruitbot environment, there remains a substantial gap in their generalization performance (as shown in Figure 6). This provides further strong evidence that DML effectively enhances the policy robustness to irrelevant features, as MDPO achieves significantly better generalization performance despite comparable training performance.
E.4 Additional Visualizations
In addition to generating adversarial samples through convolutional layers randomly initialized with a standard Gaussian distribution, we also attempt to generate adversarial samples by adjusting the brightness, contrast, saturation, and hue of the images, and test the robustness of the PPO encoder and our MDPO encoder, as shown in Figure 13.
Figure 13: The robustness of PPO and MDPO to brightness, contrast, saturation, and hue.
We can see that the MDPO policy has also learned robustness representations to these irrelevant factors, while the PPO policy remains sensitive to them. Additionally, we present adversarial samples generated by random CNNs, as shown in Figure 14, as well as those generated by randomly adjusting brightness, contrast, saturation, and hue, as can be seen from Figure 15.
Figure 14: Adversarial samples generated by random CNNs.
Figure 15: Adversarial samples generated by different brightness, contrast, saturation, and hue.
Appendix F Connections between Invariant Representation and Policy Robustness
In this section, we emphasize the key connections between invariant representation and policy robustness discussed in this paper. In the framework of invariant representation learning [29], the reinforcement learning policy $\pi$ is decoupled into
$\pi=h\circ g,$ (19)
where $g$ is usually an upstream encoder, while $h$ is an MLP that maps the latent state $z=g(o)$ ($o$ is the observation) to an action distribution. Next, we will present the precise definition of invariant representation and the policy robustness to irrelevant features introduced in this paper:
Definition F.1 (Invariant representation).
We say that the policy $\pi=h\circ g$ learns an invariant representation of high-dimensional observations, if there exists a small constant $\delta>0$, for any given underlying state $s\in\mathcal{S}$ and any two rendering functions $\tilde{f},f\in\mathcal{F}$, such that
$\left\|g(\tilde{f}(s))-g(f(s))\right\|\leq\delta.$ (20)
Definition F.2 (Policy robustness to irrelevant features).
We say that the policy $\pi$ is robust to irrelevant features of high-dimensional observations, if there exists a small constant $\xi>0$, for any given underlying state $s\in\mathcal{S}$ and any two rendering functions $\tilde{f},f\in\mathcal{F}$, such that
$\left\|\pi(\cdot|\tilde{f}(s))-\pi(\cdot|f(s))\right\|\leq\xi.$ (21)
We first assume that the downstream MLP $h$ satisfies $L$-Lipschitz continuous:
Assumption F.3.
for any given pair of latent state, $z_{1},z_{2}$, we assume that
$\left\|h(\cdot|z_{1})-h(\cdot|z_{2})\right\|\leq L\cdot\left\|z_{1}-z_{2}\right\|.$ (22)
Next, we will prove that the policy robustness to irrelevant features is more general than invariant representation. To achieve this, we first show that when the policy $\pi=h\circ g$ satisfies Definition F.1, then it also satisfies Definition F.2.
Proof.
Suppose Definition F.1 holds, according to Assumption F.3, we have
$\begin{split}\left\|\pi(\cdot|\tilde{f}(s))-\pi(\cdot|f(s))\right\|&=\left\|h(\cdot|g(\tilde{f}(s)))-h(\cdot|g(f(s)))\right\|\\ &\leq L\cdot\left\|g(\tilde{f}(s))-g(f(s))\right\|\\ &\leq L\cdot\delta,\\ \end{split}$ (23)
thus we can simply set $\xi=L\cdot\delta$, and Definition F.2 holds, concluding the proof. ∎
Then, we only need to prove that when Definition F.2 holds, Definition F.1 may not hold. To prove this, consider the following counterexample:
Proof.
Assume that the downstream MLP $h$ always outputs the same distribution over action space, then Definition F.2 holds, since
$\left\|\pi(\cdot|\tilde{f}(s))-\pi(\cdot|f(s))\right\|=\left\|h(\cdot|g(\tilde{f}(s)))-h(\cdot|g(f(s)))\right\|=0.$ (24)
In this case, it follows that the choice of $g$ is arbitrary. Therefore, Definition F.1 does not necessarily hold, concluding the proof. ∎
Building upon these proofs, we have formally shown that the policy robustness to irrelevant features (Definition F.2) is a more general definition that covers invariant representation (Definition F.1). Furthermore, the advantage of this is that we do not need to decouple the policy $\pi$ into $\pi=h\circ g$, which facilitates our theoretical analysis (Section 3).
Appendix G Proofs